OpenStudy (anonymous):

I hope to be given problems on solids of revolution and surfaces of revolution as a review for my exam =)

OpenStudy (anonymous):

i hope that your answers will be in the form of word problems. =)

OpenStudy (anonymous):

if not, then it's ok. =)

OpenStudy (cwrw238):

heres an easy one to start Find volume generated when the area between x = 0 and x = 2 under the curve y=x^2 is rotated through 360 degrees about the x-axis.

OpenStudy (anonymous):

ok. i'll draw first. which do you prefer, shells or disks? or both? =)))

OpenStudy (cwrw238):

hmmmm - disks

OpenStudy (anonymous):

via disks: \[A =\pi \int\limits_{0}^{2} (x^{2})^{2} dx = \pi \int\limits_{0}^{2} x^{4} dx\]

OpenStudy (cwrw238):

to be honest , maybe its because i'm from UK i only know the disk method

OpenStudy (cwrw238):

shell?

OpenStudy (anonymous):

32/5 square units. =))) . yes, cylindrical shells. =)))

OpenStudy (anonymous):

this problem will get too complicated when done with cylindrical shells. XD =)))

OpenStudy (cwrw238):

ok

OpenStudy (cwrw238):

hold on - i've got another for you

OpenStudy (anonymous):

answer via cylindrical shells: \[A = 2\pi \int\limits_{0}^{4} y ( 2- \sqrt{y} ) dy\]

OpenStudy (anonymous):

see, needlessly complicated. =))) for the same answer =)))

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

try this one find volume generated when the area external to the hyperbola xy = 3, in the first quadrant, bounded by the lines x=2, y = 3 is rotated about the line x=2 through 2 pi radians. a diagram would obviously help here.

OpenStudy (anonymous):

2pi radians means one whole rotation? aye?

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):



OpenStudy (cwrw238):

excuse the drawing!!!

OpenStudy (anonymous):

it's ok. good thing.; i misinterpreted it... external to parabola means over the graph or under da graph? =)))

OpenStudy (anonymous):

to the hyperbola, i mean... =))

OpenStudy (cwrw238):